# ”Security for Everyone” in Finite Blocklength IRS-aided Systems With Perfect and Imperfect CSI

### v2

@@ IV-1 Sub-Problem for Optimizing the Beamforming Vectors
Next, we show the steps to derive the LMI of equations (36e) and (36f). Using (12), the inequalities (72) and (73) in Appendix D, and by defining $x\triangleq V_{i}$, $\mathbf{A}\triangleq\rho_{i}$, and $B\triangleq\varsigma_{i}$, the dispersion factors for $i\in\{k,e\}$ can be expressed as:
Algorithm 2 PCCP Algorithm to solve problem $(\mathcal{P}2.3)$
@@
1:Initialize: $(\boldsymbol{w}^{(1)},\boldsymbol{\theta}^{(1)})$, the maximum penalty factor $o_{\text{max}}$, the scaling factor $\nu\geq 1$, convergence tolerance $\epsilon_{t_{1}}$ and $\epsilon_{t_{2}}$, maximum number of iterations $\iota_{\text{max}}$, and Set $\iota=1$.
@@
3:For $\iota\leq\iota_{\text{max}}$
4:Solve problem $(\mathcal{P}2.3)$ to obtain $\boldsymbol{\theta}^{(\iota+1)}$;
5:end if $\left\|e^{(j\boldsymbol{\theta}^{(\iota)})}-e^{(j\boldsymbol{\theta}^{(\iota-1)})}\right\|_{1}\leq\epsilon_{t_{1}}$ and $Q\leq\epsilon_{t_{2}}$.
@@
7:Obtain $o^{(\iota+1)}=\text{min}\{\nu o^{(\iota)},o_{\text{max}}\}$;
8:Update $\iota=\iota+1$;
9:else
10:define a new $\boldsymbol{\theta}^{(1)}$, set $\nu>1$ and $\iota=0$.
11:Output the feasible solution $\boldsymbol{\theta}^{*}=\boldsymbol{\theta}^{(\iota)}$.
$\xi_{i}\sqrt{V_{i}}$ $\leq\frac{\xi_{i}\sqrt{V_{i}^{(\iota)}}}{2}+\frac{\xi_{i}|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}|^{2}}{\sqrt{V_{i}^{(\iota)}}\left(\|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{k}\|_{2}^{2}+\sigma_{i}\right)},$ (55)
$\frac{(\tilde{\varphi}_{i}-\mathfrak{L})}{\mathfrak{Y}}\geq\frac{|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}|^{2}}{\|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{k}\|_{2}^{2}+\sigma_{i}},\leavevmode\nobreak\ i\in\{k,e\},$ (56)
where $\mathfrak{L}\triangleq\frac{\xi_{i}\sqrt{V_{i}^{(\iota)}}}{2}$ and $\mathfrak{Y}\triangleq\dfrac{\xi_{i}}{\sqrt{V_{i}^{(\iota)}}}$.
$\left|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k}\right|^{2}\leq\left(\frac{(\tilde{\varphi}_{i}-\mathfrak{L})}{\mathfrak{Y}}\right)\zeta_{i},\forall k,i\in\{k,e\},$ (57a)
$\left\|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{k}\right\|^{2}+\sigma_{i}\geq\zeta_{k},\forall k,i\in\{k,e\}.$ (57b)
Similar to (54), we can express (57b) by its equivalent LMI as:
$\begin{bmatrix}\widetilde{C}_{11,i}&-\tilde{c}^{H}_{i}&\mathbf{0}_{1\times M}\\ -\tilde{c}_{i}&\mathbf{I}_{K}&-\Omega_{i}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{k})^{H}\\ \mathbf{0}_{M\times 1}&-\Omega_{i}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{W}_{k})&\mathfrak{u}_{i}\mathbf{I}_{M}\end{bmatrix}\succeq 0,j\neq k.$ (58)
where $\widetilde{C}_{11,i}\triangleq\zeta_{k}-\sigma_{k}-\mathfrak{u}_{i}$, $\mathfrak{u}\triangleq[\mathfrak{u}_{1},\mathfrak{u}_{2},\dots,\mathfrak{u}_{i}]\geq 0$ is a slack variable, and $\tilde{c}_{i}\triangleq((\widehat{\boldsymbol{u}}_{i}^{H}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}))\mathbf{W}_{k})^{H}$.
Next, we can express (57a) as the following LMI:
$\begin{bmatrix}C_{k_{f}}&\tilde{c}^{H}_{ke}&\mathbf{0}_{1\times M}\\ \tilde{d}_{i}&1&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k})&\tilde{\mathfrak{u}}_{i}\mathbf{I}_{M}\end{bmatrix}\succeq 0,j\neq k.$ (59)
@@
where $C_{k_{f}}\triangleq\frac{\tilde{\varphi}_{k}\zeta_{k}}{\mathfrak{Y}}-\frac{\mathfrak{L}\zeta_{k}}{\mathfrak{Y}}-\tilde{\mathfrak{u}}_{i}$, $\tilde{\boldsymbol{\mathfrak{u}}}\triangleq[\tilde{\mathfrak{u}}_{1},\dots,\tilde{\mathfrak{u}}_{K}]\geq 0$ is a slack variable, and $\tilde{d}_{i}\triangleq((\widehat{\boldsymbol{u}}_{e}\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k})^{H}$.
@@
However, (59) is still non-convex due to the non-convex nature of the term $\tilde{\varphi}\zeta_{k}$. Similar to (44), we employ the SCA technique to tackle the non-convex expression as:
$\left(\tilde{\varphi}\zeta_{i}\right)^{ub}\triangleq\tilde{\varphi}^{(\iota)}\zeta_{i}+(\tilde{\varphi}-\tilde{\varphi}^{(\iota)})\zeta_{i}^{(\iota)}$ (60)
Algorithm 3 Proposed AO Algorithm for Solving Problem $(\mathcal{P}2)$
1:Initialize: $(\boldsymbol{w}^{(1)}=\boldsymbol{w}^{*},\boldsymbol{\theta}^{(1)}=\boldsymbol{\theta}^{*})$, $\iota=1$, and convergence tolerance $\epsilon_{t}>0$.
2:Repeat
3:Update $\boldsymbol{w}^{(\iota+1)}$ by solving $(\mathcal{P}2.2)$, and $\boldsymbol{\theta}^{(\iota+1)}$ by solving problem $(\mathcal{P}2.3)$;
4:if $\frac{|\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})-\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})|}{\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}\leq\epsilon_{t}$.
5:Then $\boldsymbol{\theta}^{(\iota)}\leftarrow\boldsymbol{\theta}^{(\iota+1)}$, $\boldsymbol{w}^{(\iota)}\leftarrow\boldsymbol{w}^{(\iota+1)}$ and terminate.
6:Otherwise $\iota\leftarrow\iota+1$ and continue.
7:Output $(\boldsymbol{w}^{*}=\boldsymbol{w}^{(\iota)}_{k},\boldsymbol{\theta}^{*}=\boldsymbol{\theta}^{(\iota)})$.
@@
where $\tilde{\varphi}^{(\iota)}$, $\zeta_{i}^{(\iota)}$ are the value of the variables $\tilde{\varphi}_{i}$, $\zeta_{i}$ at iteration $(\iota)$ in the SCA-based algorithm, respectively. Be substituting (60) in (59), we can express (59) as:
@@
$\begin{bmatrix}\widetilde{C}_{k_{f}}&\tilde{c}^{H}_{ke}&\mathbf{0}_{1\times M}\\ \tilde{d}_{i}&1&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{e}(\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}})\boldsymbol{w}_{k})&\tilde{\mathfrak{u}}_{i}\mathbf{I}_{M}\end{bmatrix}\succeq 0,j\neq k,$ (61)
@@
where $\widetilde{C}_{k_{f}}\triangleq\frac{\left(\tilde{\varphi}\zeta_{k}\right)^{ub}}{\mathfrak{Y}}-\frac{\mathfrak{L}\zeta_{k}}{\mathfrak{Y}}-\tilde{\mathfrak{u}}_{i}$.
@@
$(\mathcal{P}2.2):$ $\leavevmode\nobreak\ \underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ z$ (62a)
s.t. $z\leq{\varphi}_{k}-\mu_{k_{e}}-\tilde{\varphi}_{k}-\tilde{\varphi}_{k_{e}},\forall k,$ (62b)
@@
$\boldsymbol{\varpi}\geq 0,\boldsymbol{\varkappa}\geq 0,\boldsymbol{\vartheta}\geq 0,\boldsymbol{\varrho}\geq 0,\mathfrak{u}\geq 0,\tilde{\mathfrak{u}}\geq 0,$ (62d)
@@
At this point, all nonlinear constraints (36c), (36d), (36e), and (36f) have been linearized as LMIs (46), (51), (53), (54), (58), and (61); thus, problem ($\mathcal{P}2.2$) is an SDP problem that can be efficiently solved using standard solvers, e.g., the interior point method or the CVX toolbox [30].
@@ IV-2 Sub-Problem for Optimizing the PREs
To apply the PCCP method, we introduce an auxiliary variable set $\mathbf{b}\triangleq\{b_{n}|n\in N\}$ satisfying $b_{n}\triangleq|e^{(j\theta_{n})}|^{*}|e^{(j\theta_{n})}|$. Then, (35c) can be expressed as $b_{n}\leq|e^{(j\theta_{n})}|^{*}|e^{(j\theta_{n})}|\leq b_{n}$. The non-convex part $b_{n}\leq|e^{(j\theta_{n})}|^{*}|e^{(j\theta_{n})}|$ can be approximated by $b_{n}\leq 2\Re\{|e^{(j\theta_{n})}|^{*}|e^{(j\theta^{(\iota)}_{n})}|-|e^{(j\theta^{(\iota)}_{n})}|^{*}|e^{(j\theta^{(\iota)}_{n})}|\}$. Following the PCCP framework, we penalize the objective function (36a), hence, problem $(\mathcal{P}2.1)$ can be recast as:
$(\mathcal{P}2.3):$ $\underset{\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ z-oQ,$ (63a)
@@
$|e^{(j\theta_{n})}|^{*}|e^{(j\theta_{n})}|\leq b_{n}+c_{n},$ (63c)
$b_{n}-\hat{c}_{n}\leq 2\Re\left\{|e^{(j\theta_{n})}|^{*}|e^{(j\theta^{(\iota)}_{n})}|\right\}$
$-|e^{(j\theta^{(\iota)}_{n})}|^{*}|e^{(j\theta^{(\iota)}_{n})}|,$ (63d)
$b_{n}\geq 0,\forall n\in N,$ (63e)
where $Q\triangleq{\sum}_{n=1}^{N}c_{n}+\hat{c}_{n}$ is the penalty term, $\mathbf{c}\triangleq\{c_{n},\hat{c}_{n}\}$ is the slack variable imposed over the modulus constraint (35c), and $o$ is the regularization factor. The regularization factor is imposed to control the UMC constraint (35c) by scaling the penalty term $Q$.
Problem $(\mathcal{P}2.3)$ is an SDP problem that can be solved by the CVX toolbox. Unlike the conventional SDR method to tackle UMC, the PCCP method, summarized in Algorithm 2, is guaranteed to find a feasible point for problem $(\mathcal{P}2.3)$ [35]. The following points are helpful for the numerical implementation of the Algorithm 2:
@@
    We invoke $o_{\text{max}}$ to avoid numerical complications if $o$ grows too large[36];
@@
    The stopping criteria $Q\leq\epsilon_{t_{2}}$ guarantees the UMC (35c) when $\epsilon_{t_{2}}$ is small [37];
@@
    The convergence of Algorithm 2 is controlled by $\left\|e^{(j\boldsymbol{\theta}^{(\iota+1)})}-e^{(j\boldsymbol{\theta}^{(\iota)})}\right\|_{1}\leq\epsilon_{t_{1}}$.

### v3

@@ IV-1 Sub-Problem for Optimizing the Beamforming Vectors
Next, we show the steps to derive the LMI of equations (36e) and (36f). Using (12), the inequalities (69) and (70) in Appendix D, and by defining $x\triangleq V_{i}$, $\mathbf{A}\triangleq\rho_{i}$, and $B\triangleq\tau_{i}$, the dispersion factors for $i\in\{k,e\}$ can be expressed as:
$\xi_{i}\sqrt{V_{i}}$ $\leq\frac{\xi_{i}\sqrt{V_{i}^{(n)}}}{2}+\frac{\xi_{i}|(\boldsymbol{l}_{i}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k}|^{2}}{\sqrt{V_{i}^{(n)}}\left(\|(\boldsymbol{l}_{i}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{W}_{k}\|_{2}^{2}+\sigma_{i}\right)},$
$\frac{(\tilde{\beta}_{i}-\mathfrak{L})}{\mathfrak{Y}}\geq\frac{|(\boldsymbol{l}_{i}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k}|^{2}}{\|(\boldsymbol{l}_{i}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{W}_{k}\|_{2}^{2}+\sigma_{i}},$ (55)
where $\mathfrak{L}\triangleq{\xi_{i}\sqrt{V_{i}^{(n)}}}/{2}$ and $\mathfrak{Y}\triangleq{\xi_{i}}/{\sqrt{V_{i}^{(n)}}}$.
$\left|(\boldsymbol{l}_{i}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k}\right|^{2}\leq\zeta_{i}{(\tilde{\beta}_{i}-\mathfrak{L})}/{\mathfrak{Y}},\forall k,i\in\{k,e\},$ (56a)
$\left\|(\boldsymbol{u}_{i}\boldsymbol{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{W}_{k}\right\|^{2}+\sigma_{i}\geq\zeta_{k},\forall k,i\in\{k,e\}.$ (56b)
Similar to (54), we can express (56b) by its equivalent LMI as:
$\begin{bmatrix}\widetilde{C}_{11,i}&-\tilde{c}^{H}_{i}&\mathbf{0}_{1\times M}\\ -\tilde{c}_{i}&\mathbf{I}_{K}&-\Omega_{i}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{W}_{k})^{H}\\ \mathbf{0}_{M\times 1}&-\Omega_{i}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{W}_{k})&\mathfrak{l}_{i}\mathbf{I}_{M}\end{bmatrix}\succeq 0,$ (57)
where $\widetilde{C}_{11,i}\triangleq\zeta_{k}-\sigma_{k}-\mathfrak{u}_{i}$, $\boldsymbol{\mathfrak{l}}\triangleq[\mathfrak{l}_{1},\mathfrak{u}_{2},\dots,\mathfrak{l}_{\mathcal{K}}]\geq 0$ is a slack variable, and $\tilde{c}_{i}\triangleq((\widehat{\boldsymbol{l}}_{i}^{H}\mathbf{\Phi}\boldsymbol{L}_{\text{AR}}))\mathbf{W}_{k})^{H}$.
Next, we can express (56a) as the following LMI:
$\begin{bmatrix}D_{11,i}&\tilde{d}_{i}^{H}&\mathbf{0}_{1\times M}\\ \tilde{d}_{i}&1&\Omega_{e}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{e}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k})&\tilde{\mathfrak{l}}_{i}\mathbf{I}_{M}\end{bmatrix}\succeq 0,$ (58)
Algorithm 2 PCCP Algorithm for solving $(\mathcal{P}2.3)$
@@
1:Initialize: $(\boldsymbol{w}^{(1)},\boldsymbol{\theta}^{(1)})$, penalty factor $a_{\max}$, scaling $\nu\!\geq\!1$, tolerances $(\epsilon_{t_{1}},\epsilon_{t_{2}})$, $n_{\max}$, set $n=1$.
@@
3:  Solve $(\mathcal{P}2.3)$ to obtain $\boldsymbol{\theta}^{(n+1)}$.
4:  if $\|e^{j\boldsymbol{\theta}^{(n)}}-e^{j\boldsymbol{\theta}^{(n-1)}}\|_{1}\leq\epsilon_{t_{2}}$ and $C\leq\epsilon_{t_{1}}$ then
5:   break
@@
7:   Update $a^{(n+1)}=\min\{\nu a^{(n)},a_{\max}\}$, set n=n+1n\!=\!n+1.
8:  end if
9:until $n>n_{\max}$
10:Output: feasible solution $\boldsymbol{\theta}^{*}=\boldsymbol{\theta}^{(n)}$.
@@
where $D_{11,i}\triangleq\frac{\tilde{\beta}_{i}\zeta_{i}}{\mathfrak{Y}}-\frac{\mathfrak{L}\zeta_{i}}{\mathfrak{Y}}-\tilde{\mathfrak{l}}_{i}$, $\tilde{\mathfrak{l}}\triangleq[\tilde{\mathfrak{l}}_{1},\dots,\tilde{\mathfrak{l}}_{\mathcal{K}}]\geq 0$ is a slack variable, and $\tilde{d}_{i}\triangleq((\widehat{\boldsymbol{l}}_{e}\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k})^{H}$.
@@
However, (58) is still nonconvex due to the nonconvex nature of the term $\tilde{\beta}\zeta_{k}$. Similar to (44), we employ the SCA technique to tackle the nonconvex expression as:
$\left(\tilde{\beta}\zeta_{i}\right)^{ub}\triangleq\tilde{\beta}^{(n)}\zeta_{i}+(\tilde{\beta}-\tilde{\beta}^{(n)})\zeta_{i}^{(n)}$ (59)
@@
where $\tilde{\beta}^{(n)}$, $\zeta_{i}^{(n)}$ are the value of the variables $\tilde{\beta}_{i}$, $\zeta_{i}$ at iteration $(n)$ in the SCA-based algorithm, respectively. Be substituting (59) in (58), we can express (58) as:
@@
$\begin{bmatrix}\widetilde{D}_{i}&\tilde{b}^{H}_{ke}&\mathbf{0}_{1\times M}\\ \tilde{d}_{i}&1&\Omega_{e}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k})^{H}\\ \mathbf{0}_{M\times 1}&\Omega_{e}(\mathbf{\Phi}\boldsymbol{L}_{\text{AR}})\boldsymbol{w}_{k})&\tilde{\mathfrak{l}}_{i}\mathbf{I}_{M}\end{bmatrix}\succeq 0,j\neq k,$ (60)
@@
where $\widetilde{D}_{i}\triangleq\frac{\left(\tilde{\beta}\zeta_{i}\right)^{ub}}{\mathfrak{Y}}-\frac{\mathfrak{L}\zeta_{i}}{\mathfrak{Y}}-\tilde{\mathfrak{l}}_{i}$.
@@
$(\mathcal{P}2.2):$ $\leavevmode\nobreak\ \underset{\boldsymbol{w},\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ \mathcal{Z}$ (61a)
s.t. $\mathcal{Z}\leq{\beta}_{k}-\Upsilon_{k_{e}}-\tilde{\beta}_{k}-\tilde{\beta}_{k_{e}},\forall k,$ (61b)
@@
$\boldsymbol{\eta}\geq 0,\boldsymbol{\varkappa}\geq 0,\boldsymbol{\vartheta}\geq 0,\boldsymbol{\varrho}\geq 0,\mathfrak{l}\geq 0,\tilde{\mathfrak{l}}\geq 0,$ (61d)
@@
At this point, all nonlinear constraints (36c), (36d), (36e), and (36f) have been linearized as LMIs (46), (51), (53), (54), (57), and (60); thus, problem ($\mathcal{P}2.2$) is an SDP that can be efficiently solved using standard solvers, e.g., the interior point method or the CVX toolbox [31].
@@ IV-2 Sub-Problem for Optimizing the PREs
Algorithm 3 Proposed AO Algorithm for Solving Problem $(\mathcal{P}2)$
1:Initialize: $(\boldsymbol{w}^{(1)}=\boldsymbol{w}^{*},\leavevmode\nobreak\ \boldsymbol{\theta}^{(1)}=\boldsymbol{\theta}^{*})$, set $n=1$, and tolerance $\epsilon_{t}>0$.
2:repeat
3:  Update $\boldsymbol{w}^{(n+1)}$ by solving $(\mathcal{P}2.2)$ and $\boldsymbol{\theta}^{(n+1)}$ by solving $(\mathcal{P}2.3)$.
4:  $\Delta\!=\!\frac{|\min_{k\in\mathcal{K}}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n+1)})-\min_{k\in\mathcal{K}}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})|}{\min_{k\in\mathcal{K}}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})}$.
5:  Update $\boldsymbol{w}^{(n)}\!\leftarrow\!\boldsymbol{w}^{(n+1)},\leavevmode\nobreak\ \boldsymbol{\theta}^{(n)}\!\leftarrow\!\boldsymbol{\theta}^{(n+1)}$, and set $n\!\leftarrow\!n+1$.
6:until $\Delta\leq\epsilon_{t}$
7:Output: $(\boldsymbol{w}^{*}=\boldsymbol{w}^{(n)},\leavevmode\nobreak\ \boldsymbol{\theta}^{*}=\boldsymbol{\theta}^{(n)})$.
@@
$(\mathcal{P}2.3):$ $\underset{\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ \mathcal{Z}-a\mathcal{T},$ (62a)
[…]
$|e^{(j\theta_{m})}|^{*}|e^{(j\theta_{m})}|\leq q_{m}+t_{m},$ (62c)
[…]
$q_{m}\geq 0,\forall m\in M,$ (62e)
where $\mathcal{T}\triangleq{\sum}_{m=1}^{N}t_{m}+\hat{t}_{m}$ is the penalty term, $\mathbf{c}\triangleq\{c_{n},\hat{c}_{n}\}$ is the slack variable imposed over the modulus constraint (35c), and $a$ is the regularization factor. The regularization factor is imposed to control the UMC constraint (35c) by scaling the penalty term $Q$.
Problem $(\mathcal{P}2.3)$ is an SDP that can be solved by the CVX toolbox. Unlike the conventional SDR method to tackle UMC, the PCCP method, summarized in Algorithm 2, is guaranteed to find a feasible point for problem $(\mathcal{P}2.3)$ [36]. The following points are helpful for the numerical implementation of the Algorithm 2:
@@
    We invoke $a_{\text{max}}$ to avoid numerical complications if $a$ grows too large[37];
@@
    The stopping criteria $\mathcal{T}\leq\epsilon_{t_{1}}$ guarantees the UMC (35c) when $\epsilon_{t_{1}}$ is small [38];
@@
    The convergence of Algorithm 2 is controlled by $\left\|e^{(j\boldsymbol{\theta}^{(n+1)})}-e^{(j\boldsymbol{\theta}^{(n)})}\right\|_{1}\leq\epsilon_{t_{2}}$.